Solve the following equations:
(2b/3) - 5 = 3.........

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

By transposing -5 from LHS to RHS it becomes 5
Then,
= 2b/3 = 3 + 5
= 2b/3 = 8
Now,
Multiply both side by 3, = (2b/3) × 3= 8 × 3
= 2b = 24
And,
Divide both side by 2,
= 2b/2= 24/2
= b = 12

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
(2b/3) - 5 = 3
2b/3 = 8
2b = 24
b = 24/2
b = 12
 
Hope it helps!
Thank you ✌️